OpenStudy (anonymous):

find the circumference of a circle a( give the exact answer in terms of pi and approximate the answer by using 3.14 for pi the R= 6.3 inch

OpenStudy (anonymous):

pi*r^2=39.69pi =124.67

OpenStudy (anonymous):

=124.63*

OpenStudy (anonymous):

so what is the circumference?

OpenStudy (anonymous):

circumference = 2*pi*R=2*3.14*6.3=39.564 inch

OpenStudy (anonymous):

area of circle is pi*r^2